OpenStudy (hockeychick23):

Write the given quotient in the form a+bi. (4-5i)/7+3i Simplify your answer

OpenStudy (midhun.madhu1987):

Multiply both numerator and denominator by the conjugate of (7+3i) which is (7-3i)

OpenStudy (midhun.madhu1987):

\[(\frac{ 4-5i }{ 7+3i })(\frac{ 7-3i }{ 7-3i })\]

OpenStudy (hockeychick23):

15i^2-47i+28/-9i^2+49

OpenStudy (midhun.madhu1987):

\[\frac{ 28 - 12i - 35i + 15i ^{2} }{ 7^{2} - (3i)^{2} }\]

OpenStudy (hockeychick23):

15i^2-47i+28/49-3i^2

OpenStudy (midhun.madhu1987):

\[\frac{ 28 - 12i - 35i + 15i ^{2} }{ 7^{2} - (3i)^{2} }\]

OpenStudy (midhun.madhu1987):

This is the next step...Please combine the like terms from the numerator... also simplify the denominator as well... Do you know the value of i^2?

OpenStudy (hockeychick23):

28-47i+15i^2/49-3i^2 i^2= -1 28-47i+15(-1)/49-3(-1) 28-47i-15/52 13-47i/52